# ReWiND: Language-Guided Rewards Teach Robot Policies without New Demonstrations

### v1

@@ A.1.3 Policy Training
Specific architectural and training details are discussed per-environment in the corresponding sections Section B.2 and Section C.2. Below we talk about high level algorithmic details for policy training along with shared implementation details across environments.
@@ B.1 Simulation Setup
We also choose another 17 tasks from the MT50 benchmark for reward model evaluation and 8 of tasks are selected for downstream policy finetuning.333These 17 tasks were chosen for sharing at least some characteristic with a training task. The evaluation tasks include Window-Close, Sweep-Into, Soccer, Reach-Wall, Push-Back, Plate-Slide-Side, Plate-Slide-Back, Pick-Place-Wall, Handle-Pull, Handle-Press-Side, Faucet-Close, Door-Lock, Door-Close, Coffee-Push, Coffee-Button, Button-Press-Wall, Button-Press-Topdown. These tasks are visually similar to the training tasks, but the tasks are different. The 8 tasks used for downstream policy training are Window-Close, Reach-Wall, Handle-Pull, Coffee-Button, Button-Press-Wall, Door-Lock, Handle-Press-Side, Sweep-into.
@@ B.2 Training Details
We pre-train the policy with IQL [51] for 100K steps with learning rate 0.001, gamma 0.99. We use a three layer MLP of size for both the policy and value function network. The general training procedure is described in (Section A.1.3)
@@ C.1 Robot Experiment Setup
We use the Koch1.1 bimanual arm setup for data collection and learning [65].444https://github.com/jess-moss/koch-v1-1 Altogether, four total arms (2 for data collection) cost $1000, letting us demonstrate ReWiND enables real-world online RL of new tasks even with very low-cost hardware and noisy control. The observations consist of RGB images from a Logitech C930e top camera and side camera (pictured in Figure 11). We control the robot with absolute joint position control at a frequency of 30Hz. We collect a small dataset of 10 demonstrations over 20 tasks, and then use 5 demos per-task for the reward function. We found the offline-trained policy to be the primary bottleneck to optimizing rewards in unseen tasks, so we used 10 demos per-task for offline policy training. We have an episode timeout of 250 steps and provide a success bonus of 125 upon success (from Equation 4 and Equation 5). Proprioceptive information in this environment includes 12 robot joint states, 6 for each arm. These represent the rotation of each joint and gripper..
@@ C.2 Real Robot Training Details
Unlike the MetaWorld experiments that use an MLP-based policy, we use an action-chunked policy with temporal ensembling for the real robot. We found chunking to lead to more stable bimanual manipulation on the Koch arms. We implement the action chunking with a Transformer policy that predicts 60 actions at each timesteps corresponding to 2 seconds of actions. We also implement a Transformer-based critic. During rollouts, we then use temporal ensembling [91]. Here, the current action is ensembles with the last 60 timesteps’ predictions according to an exponential weighting scheme , where we use or depending on the task. We found to work well for tasks requiring grasping solid objects as it weights recent actions more heavily, necessary for ensuring the policy actually commits to the grasp, and to work well for non-grasping tasks as it results in a smoother policy.
@@
We set in both equations to a fixed value of 10.0 on tasks where grasping solid objects is not required. For others, we set it to 20.0 to ensure the policy doesn’t degenerate from its grasping action early in training. We found that lower KL penalties could result in the policy falling into locally optimal but globally suboptimal behaviors, such as moving a cup with the arm instead of actually picking it up.
@@ C.3 Real Robot Tasks
We collected 10 demos per-task over 20 tasks on the Koch arms. We train the reward function on 5 demos per-task and the policy on 10 demos per-task. We list these training tasks below.

### v2

@@ A.1.3 Policy Training
Specific architectural and training details are discussed per-environment in the corresponding sections Section B.2 and Section C.2. Below we talk about high-level algorithmic details for policy training along with shared implementation details across environments.
@@ B.1 Simulation Setup
We also choose another 17 tasks from the MT50 benchmark for reward model evaluation and 8 of tasks are selected for downstream policy finetuning.333These 17 tasks were chosen for sharing at least some characteristic with a training task. The evaluation tasks include Window-Close, Sweep-Into, Soccer, Reach-Wall, Push-Back, Plate-Slide-Side, Plate-Slide-Back, Pick-Place-Wall, Handle-Pull, Handle-Press-Side, Faucet-Close, Door-Lock, Door-Close, Coffee-Push, Coffee-Button, Button-Press-Wall, Button-Press-Topdown. We select these training tasks and target evaluation tasks to test directional and spatial (e.g., door open → close, button press → + wall) generalization scenarios where the policy is unlikely to be able to solve the task, due to requiring significantly different action sequences from the most visually similar training task, while a good reward function should be able to guide the policy to solve it. The 8 tasks used for downstream policy training are Window-Close, Reach-Wall, Faucet-Close, Coffee-Button, Button-Press-Wall, Door-Lock, Handle-Press-Side, Sweep-into.
@@ B.2 Training Details
We pre-train the policy with IQL [52] for 100K steps with learning rate 0.001, discount . We use a three-layer MLP of size for both the policy and value function network. The general training procedure is described in Section A.1.3
@@ C.1 Robot Experiment Setup
We use the Koch1.1 bimanual arm setup for data collection and learning [66].444https://github.com/jess-moss/koch-v1-1 Altogether, four total arms (2 for data collection) cost $1000, letting us demonstrate ReWiND enables real-world online RL of new tasks even with very low-cost hardware and noisy control. The observations consist of RGB images from a Logitech C930e top camera and side camera (pictured in Figure 12). We control the robot with absolute joint position control at a frequency of 30Hz. We collect a small dataset of 10 demonstrations over 20 tasks, and then use 5 demos per-task for the reward function. We found the offline-trained policy to be the primary bottleneck to optimizing rewards in unseen tasks, so we used 10 demos per-task for offline policy training. We have an episode timeout of 250 steps and provide a success bonus of 125 upon success (from Equation 4 and Equation 5). Proprioceptive information in this environment includes 12 robot joint states, 6 for each arm. These represent the rotation of each joint and gripper.
@@ C.2 Real Robot Training Details
Unlike the MetaWorld experiments that use an MLP-based policy, we use an action-chunked policy with temporal ensembling for the real robot. We found chunking to lead to more stable bimanual manipulation on the Koch arms. We implement the action chunking with a Transformer policy that predicts 60 actions at each timestep corresponding to 2 seconds of actions. We also implement a Transformer-based critic. During rollouts, we then use temporal ensembling [99]. Here, the current action is ensembled with the last 60 timesteps’ predictions according to an exponential weighting scheme , where we use or depending on the task. We found to work well for tasks requiring grasping solid objects as it weights recent actions more heavily, necessary for ensuring the policy actually commits to the grasp, and to work well for non-grasping tasks as it results in a smoother policy.
@@
We set in both equations to a fixed value of 10.0 on tasks where grasping solid objects is not required. For others, we set it to 20.0 to ensure the policy doesn’t degenerate from its grasping action early in training. We found that lower KL penalties could result in the policy falling into locally optimal but globally suboptimal behaviors, such as moving a cup with the arm instead of actually picking it up. Work released after our paper further explores RL in chunked action spaces with similar KL-regularized behavior constraints, and find that such constraints and chunking helps improve exploration during RL [101]
@@ C.3 Real Robot Tasks
We collected 10 demos per-task over 20 tasks on the Koch arms. We train the reward function on 5 demos per-task and the policy on 10 demos per-task. We selected these training and evaluation tasks specifically to evaluate spatial, visual, and language generalization. We list these training tasks below.